OpenStudy (steve816):

Please I need your help! Find the average rate of change from 2 to 3 for each function f.

OpenStudy (steve816):

\[f(x)=3x-4x^2\]

OpenStudy (anonymous):

\[average. rate =\frac{ f(b)-f(a) }{ b-a }\]

OpenStudy (anonymous):

so \[f(x=2)=3*2-4(2)^2=6-16=-10\]

OpenStudy (anonymous):

similarly \[f(x=3)=-27\]

OpenStudy (anonymous):

\[average rate=\frac{ -27-(-10) }{ 3-2 }=-17\]

OpenStudy (anonymous):

now if it was a practical question, we would have units -17 units per unit time